solve the equation for y.
4x+6y=24

Answer by srinivas.g(540)

4x +6y = 24
subtrat 4x on both sides
4x+6y-4x = 24-4x
6y =24-4x
divide with 6 on both sides
6y/6 = (24-4x)/6
y= 24/6 - 4x/6
y= 4 - 2/3 x